# MNO: Multiscale Neural Operator for Computational Fluid Dynamics with 3D Point Cloud Data

### v1

@@ Abstract
Neural operators have emerged as a powerful data-driven paradigm for solving Partial Differential Equations (PDEs), offering orders-of-magnitude acceleration over traditional solvers. However, existing approaches still suffer from limited accuracy and scalability, particularly on irregular domains where fluid flows exhibit rich multiscale structures. In this work, we introduce the Multiscale Neural Operator (MNO), a new architecture for Computational Fluid Dynamics (CFD) on three-dimensional (3D) unstructured point clouds. MNO explicitly decomposes information across three scales: a global dimension-shrinkage attention module for long-range dependencies, a local graph attention module for neighborhood-level interactions, and a micro point-wise attention module for fine-grained details. This design preserves multiscale inductive biases while remaining computationally efficient. We evaluate MNO on four diverse benchmarks, covering both steady-state and unsteady flow scenarios with up to 300K points. Across all tasks, MNO consistently outperforms state-of-the-art baselines, reducing prediction errors by 5% to 40% and demonstrating improved robustness in challenging 3D CFD problems. Our results highlight the importance of explicit multiscale design for neural operators and establish MNO as a scalable framework for learning complex fluid dynamics on irregular domains.
@@ 1 Introduction
Despite this remarkable efficiency, neural operators still fall short of traditional solvers in accuracy (typically $10^{-3}$ versus $10^{-7}$ relative error). Recent works have sought to close this gap through carefully‑designed feature transformations, including spectral mappings FNO (Li et al., 2021), global latent space learning LNO (Wang & Wang, 2024), and Transformer-stacked approach with multiple physical space transformations Transolver (Wu et al., 2024), etc. Yet the intrinsic multiscale nature of fluid flow remains (Rahman et al., 2023; Wen et al., 2022) largely underexplored in architectural design, particularly on irregular and unstructured domains. In addition, sole reliance on global modeling often sacrifices local details, while fine-grained attention mechanisms incur prohibitive computational costs. These challenges highlight the need for architectures that explicitly disentangle and integrate information across multiple spatial scales.
@@
In this work, we propose a Multiscale Neural Operator (MNO) to tackle typical CFD tasks on irregular domains. The motivation stems from the observation that physical quantities in flow fields exhibit strong multiscale effects: large-scale global trends, localized interactions near object surfaces, and fine-grained pointwise variations. Our goal is to develop a general framework that can faithfully represent objects in three-dimensional (3D) flow fields and accurately predict critical physical quantities such as pressure and velocity.
@@
At the core of MNO is a sequence of three-scale blocks, each combining three complementary, parallel modules: (1) a Global Dimension-Shrinkage Attention module, which projects $N$ points into a compact set of $M$ modes to capture long-range dependencies; (2) a Local Graph Attention module, which encodes $k$-nearest-neighbor interactions to model mid-scale neighborhood dynamics; and (3) a Micro Point-wise Attention module, which evolves each point’s features independently to retain high-frequency variations. The outputs of these modules are fused after each block, enabling MNO to integrate receptive fields across scales and capture a broad spectrum of physical phenomena. Built directly on 3D point clouds, this design avoids mesh constraints and provides a unified framework for extracting global, local, and fine-grained flow representations.
@@
We validate MNO across multiple benchmarks, spanning both steady-state and unsteady flow tasks with point resolutions ranging from 15K to 300K. Compared to state-of-the-art methods, MNO reduces prediction errors by 5% to 40%, demonstrating consistent improvements in accuracy and robustness on challenging 3D CFD problems.
In summary, the main contributions of this paper are as follows:
  • We propose a Multiscale Neural Operator (MNO) for CFD on unstructured point clouds. Unlike prior multiscale methods restricted to regular grids, MNO directly processes 3D point data, removing mesh constraints and enabling flexible modeling of complex geometries and dynamic domains.
  • We introduce an explicit multiscale decomposition with three complementary modules: global dimension-reduction attention for long-range dependencies, local graph attention for neighborhood interactions, and micro point-wise attention for fine-grained details, ensuring balanced representation across scales.
@@
  • We evaluate MNO on four diverse datasets, covering both steady and unsteady CFD tasks, and show that it consistently outperforms state-of-the-art baselines, reducing prediction errors by 5% to 40%.
@@ 2 Related Work
Deep learning for PDEs has mainly progressed along two paths: physics-informed networks that enforce PDE constraints during training, and neural operators that learn solution mappings directly from data. We briefly review both directions, emphasizing their use in fluid dynamics and their limitations in modeling multiscale predictions on irregular domains.
@@ 2.1 Physics-Informed Neural Networks
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) embed PDE constraints into the loss function, enabling solution learning without labeled data. Despite inspiring many extensions (Wang et al., 2021; 2022; Karlbauer et al., 2022; Rao et al., 2023), PINNs require task-specific loss design, struggle with unstructured point clouds, and is hard to scale to high-dimensional or stiff PDEs, limiting their applicability to complex CFD tasks.
@@ 2.2 Neural Operators
Irregular Domain Neural Operators
PointNet (Qi et al., 2017a) and PointNet++ (Qi et al., 2017b) introduce point-based learning with global pooling and hierarchical neighbor search, respectively, though the latter often incurs high cost and may lose fine-scale details. Geo-FNO (Li et al., 2023a) maps irregular meshes into a uniform latent space for FFT-based FNO operations. LNO (Wang & Wang, 2024) encodes point clouds into compact latent tokens and applies Transformer layers for global modeling, while Transolver (Wu et al., 2024) compresses tokens into physical slices for Transformer-based feature extraction. PCNO (Zeng et al., 2025) combines FNO-style global features with residual and gradient-based local features. Despite these advances, most irregular-domain operators emphasize global features, paying limited attention to the coupling between local and global scales.
Multiscale Neural Operators
U-NO (Rahman et al., 2023) integrates U-Net with neural operators, enabling multiscale PDE mapping. U-FNO (Wen et al., 2022) enhances FNO with local convolutions, while MscaleFNO (You et al., 2024) employs multiple FNO branches to extract features at different scales. These methods improve multiscale representation but remain tied to regular grids, limiting their ability to handle geometric deformations and dynamic flow fields. This motivates the development of multiscale neural operators tailored to point clouds, where scale interactions can be explicitly modeled in unstructured domains.
@@ 3 Method
Figure 1: (a) The overview of the proposed MNO model with a sequence of three-scale blocks, and (b) each block combines three complementary, parallel modules: (c) a global dimension-shrinkage attention module for long-range dependencies, (d) a local graph attention module for neighborhood-level interactions, and (e) a micro point-wise attention module for fine-grained details.
@@ 3.1 Overview of the MNO model
As a concrete example, in the ShapeNet Car benchmark, after preprocessing (Deng et al., 2024; Wu et al., 2024), the input consists of $N$ points with 3D coordinates $pos_{in}$ and features $feature_{in}$ that include surface normals and signed distance values (Euclidean distance from each air point to the nearest surface point, positive outside the car). This results in an input dimension of $\mathbb{R}^{N\times 7}$. The output $X_{out}$ includes the velocity vector field in the air region and the pressure scalar field on the car surface, with an output dimension of $\mathbb{R}^{N\times 4}$.
@@ 3.2 Global Dimension-Shrinkage Attention Module
The global module captures long-range dependencies in the point cloud, enabling the model to extract global patterns such as overall shape and large-scale flow trends.
To force this module capture long-range, low-frequency feature, and to address the quadratic computational cost of applying attention on all tokens, we introduce a low rank projection strategy similar to Transolver (Wu et al., 2024) and LNO (Wang & Wang, 2024). Specifically, the latent features $X\in\mathbb{R}^{N\times D}$ are projected into a compact $M$-dimensional subspace ($M\ll N$) using a learnable projector $P$, then the feature is processed by a multi-head self-attention (MSA) in the reduced space $\mathbb{R}^{M\times D}$, and finally recover to the point feature space $\mathbb{R}^{N\times D}$ via the inverse projection $Q$. Formally, the global feature $X_{global}\in\mathbb{R}^{N\times D}$ is computed by
$\begin{split}P&=\text{Softmax}_{N}(\text{MLP}(X)),\\ Q&=\text{Softmax}_{M}(\text{MLP}(X)),\\ X_{global}&=Q\cdot Z_{lr}=Q\cdot\text{MSA}(P^{T}\cdot X),\end{split}$ (3)
where $P,Q\in\mathbb{R}^{N\times M}$ are the projection and inverse-projection matrices with $M=256$ by default, and $\text{Softmax}_{N}(\cdot)$ and $\text{Softmax}_{M}(\cdot)$ denote the Softmax function along the $N$ and $M$ dimensions, respectively.
This mechanism removes redundant information while preserving low-frequency global components. In the reduced space, attention weights can be computed efficiently at complexity $O(M^{2}D)$ instead of $O(N^{2}D)$, and the overall cost is dominated by the projection step $O(MND)$.
@@ 3.3 Local Graph Attention Module
The Local Attention module is designed to restrict interactions to geographically nearby points, ensuring that local geometric structures are explicitly preserved. Specifically, a $k$-nearest neighbor (k-NN) graph is first constructed using the Euclidean distance of the input 3D coordinates. Each spatial point serves as a graph node, and its $k$ nearest neighbors define the local connectivity.
@@
Inspired by the Point Transformer (Zhao et al., 2021), originally developed for point cloud segmentation, the Local Graph Attention computes neighborhood features for each node by attending only to its $k$ nearest neighbors. The structure of the Local Attention module is illustrated in Figure 1(d). The local features between the node and its neighboring nodes is computed following
@@
$\begin{split}X_{local}=\text{Sum}\big(\text{Softmax}\big(\text{MLP}(Q-K_{nbr}+pos_{rel}))\odot(V_{nbr}+pos_{rel}\big)\big),\end{split}$ (4)
where $Q\in\mathbb{R}^{N\times k\times D}$ denotes the replicated features of the center node ($k$ identical features vectors); $K_{nbr},V_{nbr}\in\mathbb{R}^{N\times k\times D}$ are the features of neighboring nodes, and $pos_{rel}\in\mathbb{R}^{N\times k\times D}$ encodes relative positional offsets. The symbol $\odot$ indicates element-wise multiplication. The similarity kernel is parameterized by an MLP, and the weighted neighbor features are aggregated by summation along the dimension $k$, yielding $X_{local}\in\mathbb{R}^{N\times D}$. The computational complexity is dominated by matrix multiplication in the MLP, yielding $O(NkD^{2})$.
This formulation enforces that only spatially adjacent tokens interact directly, embedding locality into the feature learning process, which ensures that fine-scale geometric and physical properties are preserved across deeper layers. In contrast, the Global Attention module complements this design by capturing long-range dependencies, and together they enable MNO to achieve both local accuracy and global coherence in flow prediction.

### v2

# MNO: Multiscale Neural Operator for 3D Computational Fluid Dynamics

@@ Abstract
Neural operators have emerged as a powerful data-driven paradigm for solving partial differential equations (PDEs), while their accuracy and scalability are still limited, particularly on irregular domains where fluid flows exhibit rich multiscale structures. In this work, we introduce the Multiscale Neural Operator (MNO), a new architecture for computational fluid dynamics (CFD) on 3D unstructured point clouds. MNO explicitly decomposes information across three scales: a global dimension-shrinkage attention module for long-range dependencies, a local graph attention module for neighborhood-level interactions, and a micro point-wise attention module for fine-grained details. This design preserves multiscale inductive biases while remaining computationally efficient. We evaluate MNO on diverse benchmarks, covering steady-state and unsteady flow scenarios with up to 300k points. Across all tasks, MNO consistently outperforms state-of-the-art baselines, reducing prediction errors by 5% to 50%. The results highlight the importance of explicit multiscale design for neural operators and establish MNO as a scalable framework for learning complex fluid dynamics on irregular domains.
Physics-informed nerual networks, nerual operator, PDE, CFD
@@ 1 Introduction
Despite this remarkable efficiency, neural operators still fall short of traditional solvers in accuracy (typically $10^{-3}$ versus $10^{-7}$ relative error). Recent works have sought to close this gap through carefully‑designed feature transformations, including spectral mappings FNO (Li et al., 2021), global latent space learning LNO (Wang and Wang, 2024), and Transformer-stacked approach Transolver (Wu et al., 2024), etc. Yet the intrinsic multiscale nature of fluid flow remains (Rahman et al., 2023; Wen et al., 2022) largely underexplored in architectural design, particularly on irregular and unstructured domains. In addition, sole reliance on global modeling often sacrifices local details, while fine-grained attention mechanisms incur prohibitive computational costs, or lead out-of-memory even in a high-end GPU. These challenges highlight the need for architectures that explicitly disentangle and integrate information across multiple spatial scales with a balanced computational budgets.
@@
In this work, we propose a Multiscale Neural Operator (MNO) to tackle typical CFD tasks on irregular domains. The motivation stems from the observation that physical quantities in flow fields exhibit strong multiscale effects: large-scale global trends, localized interactions near object surfaces, and fine-grained pointwise variations. Our goal is to develop a general framework that can faithfully represent objects in 3D flow fields and accurately predict critical physical quantities such as pressure and velocity.
@@
At the core of MNO is a sequence of three-scale blocks, each combining three complementary, parallel modules: (1) a Global Dimension-Shrinkage Attention module, which projects $N$ points into a compact set of $M$ modes to capture long-range dependencies; (2) a Local Graph Attention module, which adopts a globally shared graph structure together with differential attention to encode $k$-nearest-neighbor interactions, thereby modeling mid-scale neighborhood dynamics; and (3) a Micro Point-wise Attention module, which evolves each point’s features independently to retain high-frequency variations. The outputs of these modules are fused after each block, enabling MNO to integrate receptive fields across scales and capture a broad spectrum of physical phenomena. Built directly on 3D point clouds, this design avoids mesh constraints and provides a unified framework for extracting global, local, and fine-grained flow representations.
@@
We validate MNO across multiple benchmarks, including wind-tunnel, magnetohydrodynamics, and parachute simulations, covering both steady and unsteady flow regimes with point resolutions ranging from 15K to 300K. Compared to state-of-the-art (SOTA) methods, MNO reduces prediction errors by 5% to 50%, demonstrating consistent improvements in accuracy and robustness on challenging 3D CFD problems. Notably, when modeling complex turbulent spatiotemporal fields, for example Magnetohydrodynamics (MHD), the proposed model leads to accuracy improvements of up to 50%.
In summary, the main contributions are as follows:
  • We propose a Multiscale Neural Operator (MNO) that extracts multiscale features directly on the full point cloud graph, rather than through hierarchical resampling. Cross-scale features are tightly coupled within each network block, enabling multiscale interaction at every stage.
  • Within each block, MNO introduces three complementary scale-specific operators: a low-rank global operator for long-range dependencies, a differential attention operator for neighborhood-scale interactions, and a point-wise weighted operator for high-frequency local variations.
@@
  • MNO is evaluated on five diverse datasets, covering both steady and unsteady CFD and MHD tasks, and show that it consistently outperforms SOTA baselines. In challenging 3D turbulent spatiotemporal modeling scenarios of magnetohydrodynamics, it achieves up to a 50% performance improvement.
@@ 2 Related Work
Deep learning for PDEs has progressed along two paths: physics-informed networks that enforce PDE constraints during training, and neural operators that learn solution mappings directly from data. We briefly review both directions, emphasizing their use in fluid dynamics and their limitations in multiscale predictions on irregular domains.
@@ 2.1 Physics-Informed Neural Networks
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) embed PDE constraints into the loss function, enabling solution learning without labeled data. Despite inspiring many extensions (Wang et al., 2021, 2022; Karlbauer et al., 2022; Rao et al., 2023), PINNs require task-specific loss design, struggle with unstructured point clouds, and is hard to scale to high-dimensional or stiff PDEs, limiting their applicability to complex CFD tasks.
@@ 2.2 Neural Operators
Irregular Domain Neural Operators PointNet (Qi et al., 2017a) and PointNet++ (Qi et al., 2017b) introduce point-based learning with global pooling and hierarchical neighbor search, respectively, though the latter often incurs high cost and may lose fine-scale details. Geo-FNO (Li et al., 2023a) maps irregular meshes into a uniform latent space for FFT-based FNO operations. LNO (Wang and Wang, 2024) encodes point clouds into compact latent tokens and applies Transformer layers for global modeling, while Transolver (Wu et al., 2024) compresses tokens into physical slices for Transformer-based feature extraction. PCNO (Zeng et al., 2025b) combines FNO-style global features with residual and gradient-based local features. HAMLET (Bryutkin et al., 2024) employs a modular input encoder to construct graph transformers for solving PDEs. PhyMPGN (Zeng et al., 2025a) leverages a learnable Laplacian to guide GNNs toward learning within the physically feasible solution space. Despite these advances, most irregular-domain operators emphasize global features, paying limited attention to the coupling between local and global scales.
Figure 1: (a) The overview of the proposed MNO model with a sequence of three-scale blocks, and (b) A MNO block combining three complementary, parallel modules: (c) a global dimension-shrinkage attention module for long-range dependencies, (d) a local graph attention module for neighborhood-level interactions, and (e) a micro point-wise attention module for fine-grained details.
Multiscale Neural Operators U-NO (Rahman et al., 2023) integrates U-Net with neural operators, enabling multiscale PDE mapping. U-FNO (Wen et al., 2022) enhances FNO with local convolutions, while MscaleFNO (You et al., 2024) employs multiple FNO branches to extract features at different scales. These methods improve multiscale representation but remain tied to regular grids, limiting their ability to handle geometric deformations and dynamic flow fields. MGKN (Li et al., 2020) and AMG (Li et al., 2025) achieve multi-scale receptive fields via stage-wise downsampling on irregular geometries. However such downsampling discards important geometric and boundary information. In contrast, MNO builds its multi-scale hierarchy on the full flow-field point cloud, with cross-scale features internally coupled at every network stage.
@@ 3.1 Overview of the MNO model
As a concrete example, in the ShapeNet-Car benchmark, after preprocessing (Deng et al., 2024; Wu et al., 2024), the input consists of $N$ points with 3D coordinates $pos_{in}$ and features $feature_{in}$ that include surface normals and signed distance values (Euclidean distance from each air point to the nearest surface point, positive outside the car). This results in an input dimension of $\mathbb{R}^{N\times 7}$. The output $X_{out}$ includes the velocity vector field in the air region and the pressure scalar field on the car surface, with an output dimension of $\mathbb{R}^{N\times 4}$.
@@ 3.2 Global Dimension-Shrinkage Attention Module
The global module captures long-range dependencies in the 3D discrete domain, enabling the model to extract global patterns such as overall shape and large-scale flow trends.
To encourage this module to capture long-range, low-frequency features, and to address the quadratic computational cost of applying attention on all tokens, we introduce a low rank projection strategy inspired by LNO (Wang and Wang, 2024).
Specifically, the latent features $X\in\mathbb{R}^{N\times D}$ are projected into a compact $M$-dimensional subspace ($M\ll N$) using a learnable projector $P$, then the feature is processed by a multi-head self-attention (MSA) in the reduced space $\mathbb{R}^{M\times D}$, and finally mapped back to the point feature space $\mathbb{R}^{N\times D}$ via the learned inverse projection $Q$. Formally, the global feature $X_{global}\in\mathbb{R}^{N\times D}$ is computed by
$\begin{split}W&=\text{MLP}(X),\ W\in\mathbb{R}^{N\times M},\\ P&=\text{Softmax}_{N}(W),\ P\in\mathbb{R}^{N\times M},\\ Q&=\text{Softmax}_{M}(W),\ Q\in\mathbb{R}^{N\times M},\\ Z_{lr}&=P^{T}\cdot X,Z_{lr}\in\mathbb{R}^{M\times D},\\ Z_{lr}^{\prime}&=\text{MSA}(Z_{lr}),\ Z_{lr}^{\prime}\in\mathbb{R}^{M\times D},\\ X_{global}&=Q\cdot Z_{lr},\ X_{global}\in\mathbb{R}^{N\times D},\end{split}$ (3)
where the matrix $W$ represents the shared weight for projection $P$ and inverse-projection $Q$ matrices with $M=256$ by default. $\text{Softmax}_{N}(\cdot)$ and $\text{Softmax}_{M}(\cdot)$ denote the Softmax function along the $N$ and $M$ dimensions, respectively. $Z_{lr}$ is the low rank space projection of $X$; $Z_{lr}^{\prime}$ denotes the global feature in the low rank space. The $X_{global}$ is the global feature in 3D space. The $P$ and the $Q$ share a weight matrix, ensuring consistent representation between low rank projection and its inverse while reducing computational cost.
This module has a global receptive field via low rank projections to learn global, slowly-varying features. In the reduced space, attention weights can be computed efficiently at complexity $O(M^{2}D)$ instead of $O(N^{2}D)$, and the overall cost is dominated by the projection step $O(MND)$. Unlike LNO, which employs a single shared low rank projection throughout the network, MNO introduces an independent low rank projection in each block. This allows multiple low-rank mappings to progressively learn different global feature subspaces layer by layer, enabling different blocks to progressively explore complementary global feature subspaces.
@@ 3.3 Local Graph Attention Module
The Local Attention module is designed to restrict interactions to geographically nearby points, ensuring that local geometric structures are explicitly preserved. Specifically, a $k$-nearest neighbor (kNN) graph is first constructed using the Euclidean distance of the input 3D coordinates. Each spatial point serves as a graph node, and its $k$ nearest neighbors define the local connectivity.
@@
Inspired by the Point Transformer (Zhao et al., 2021), originally developed for point cloud segmentation, the Local Graph Attention computes neighborhood features for each node by attending only to its $k$ nearest neighbors. The structure of the Local Attention module is illustrated in Figure 1 (d). The local features between the node and its neighboring nodes is computed following.
$\begin{split}G_{knn}&=\text{kNN}(pos_{in}),\ G_{knn}\in\mathbb{R}^{N\times k},\\ Q,K,V&=\text{MLP}(X),\ Q,K,V\in\mathbb{R}^{N\times D},\\ K_{nbr}&=\text{NNI}(G_{knn},K),\ K_{nbr}\in\mathbb{R}^{N\times k\times D},\\ V_{nbr}&=\text{NNI}(G_{knn},V),\ V_{nbr}\in\mathbb{R}^{N\times k\times D},\\ P_{nbr}&=\text{NNI}(G_{knn},pos_{in}),\ P_{nbr}\in\mathbb{R}^{N\times k\times 3},\\ X_{local}&=\text{DA}(Q_{ctr},K_{nbr},V_{nbr},P_{ctr},P_{nbr}),\end{split}$ (4)
where the function kNN stands for k-nearest neighbor; $pos_{in}$ represents the coordinate position of all nodes; $G_{knn}$ is the shared kNN graph, and NNI denotes the nearest neighbor index based on $G_{knn}$. The $Q_{ctr}=Q$ denotes the features of the center nodes; $K_{nbr}$ and $V_{nbr}$ are the features of neighboring nodes; $P_{ctr}=pos_{in}$ and $P_{nbr}$ represent the original spatial positions of the central nodes and the neighboring nodes, respectively. The function DA($\cdot$) represents the differential attention mechanism.
Unlike prior graph-based approaches (Li et al., 2020, 2025) that reconstruct neighborhood graphs across layers, we share a single kNN graph across all MNO blocks, eliminating repeated neighbor searches and reducing computational overhead. Compared to fixed-radius, kNN enforces a constant neighborhood size, providing predictable memory usage and improved robustness to density variations.
The local neighborhood interactions are captured via differential attention, which is defined as follows.
@@
$\begin{split}X_{local}&=\text{DA}(Q_{ctr},K_{nbr},V_{nbr},P_{ctr},P_{nbr})\\ &=\sum_{j=1}^{k}\alpha_{j}\odot(v_{nbr}^{j}+\Delta p^{j}),\\ \alpha_{j}&=\frac{exp(\phi_{(}\Delta x^{j}+\Delta p^{j}))}{\sum_{i=1}^{k}exp(\phi_{(}\Delta x^{i}+\Delta p^{i}))},\\ \Delta x^{j}&=q_{ctr}-k_{nbr}^{j},\ \Delta x^{j}\in\mathbb{R}^{N\times D},\\ \Delta p^{j}&=\phi_{p}(p_{ctr}-p_{nbr}^{j}),\ \Delta p^{j}\in\mathbb{R}^{N\times D},\end{split}$ (5)
where the $X_{local}\in\mathbb{R}^{N\times D}$ is the output local features; $j\in[0,k-1]$ denotes the $j$-th neighboring node within the local neighborhood; $\alpha_{j}\in\mathbb{R}^{N\times D}$ represents the channel-wise attention weight applied to each neighboring node; $\Delta x^{j}$ denotes the differential feature between the neighboring node and the center node, and $\Delta p^{j}$ represents the position-enhanced relative displacement between the neighboring node and the center node; $\phi(\cdot)$ denotes an MLP-based similarity relation mapping kernel, while $\phi_{p}(\cdot)$ represents an explicit spatial displacement mapping implemented by an MLP. The operator $\odot$ denotes element-wise multiplication, and the weighted neighboring features are aggregated by summation along the neighborhood dimension $k$. The overall computational complexity is mainly dominated by the matrix multiplications in the MLPs, resulting in a complexity of $O(NkD^{2})$.
For local feature modeling, existing methods primarily rely on discrete convolutions (Raonic et al., 2023) or dot-product self-attention (Li et al., 2025). In contrast, we model local interactions via feature differences between a center node and its neighbors, and parameterize the similarity function using a learnable MLP. Compared to dot-product similarity, feature differences offer greater robustness to spatial shifts and more directly capture local structural variations. The relative positional encodings are further incorporated to explicitly encode geometric relationships within local neighborhoods.